OpenStudy (anonymous):

Is -24-3^2 over 11 = 3 or - 273/11

OpenStudy (anonymous):

-3

zepdrix (zepdrix):

\[\frac{ -24-3^2 }{ 11 }\]Problem look like this?

OpenStudy (anonymous):

yup

zepdrix (zepdrix):

\[\frac{ -24-9 }{ 11 }=\frac{ -33 }{ 11 }= -3\] Yes, good job Amy! :)

OpenStudy (anonymous):

awseome thx :) finally got one right!

zepdrix (zepdrix):

I suppose you also deserve a medal for this one! :) hehe